OpenStudy (anonymous):

HELP PLEASE ! Solve: 3(x + 4) greater than or equal to x - 2

OpenStudy (freckles):

3(x+4)=what using distributive property

OpenStudy (anonymous):

3x + 12

OpenStudy (freckles):

so you have \[3x+12 \ge x-2 \]

OpenStudy (anonymous):

yup

OpenStudy (freckles):

get your x's together and your non x terms on the opposing side by adding and subtracting on both sides

OpenStudy (anonymous):

2x > -14

OpenStudy (anonymous):

4x is bigger or equal to -14

OpenStudy (anonymous):

It's 2x , 3x - x

OpenStudy (freckles):

\[3x+12 \ge x-2 \\ \text{ subtract } x \text{ on both sides } \\ 3x-x+12 \ge x-x-2 \\ 2x+12 \ge -2 \] now try again to subtract your 12 on both sides and you have \[2x \ge -14\]

OpenStudy (anonymous):

yup

OpenStudy (freckles):

now the last step is to isolate the x factor on the left hand side

OpenStudy (anonymous):

so it's x is bigger or equal to -7

OpenStudy (freckles):

yeah

OpenStudy (freckles):

good job

OpenStudy (anonymous):

thanks

OpenStudy (freckles):

np :)